OpenStudy (anonymous):

Find the surface area of a sphere with a radius of 1. Use 3.14 for π. Round your answer to the nearest tenth.

OpenStudy (anonymous):

surface area of a sphere is given by the formula A=4*pi*r^2 so the surface area of sphere with radius of 1 is A=4*3.14*1=12.56=12.6 (rounded to tenth)

OpenStudy (anonymous):

thhanks cuz i been stuck on that for a while.